Derivative of x^3+4xy^2+5y^3=1 at (2,-1)


1
Expert's answer
2021-10-29T02:52:11-0400

Differentiate both sides with respect to "x"


"\dfrac{d}{dx}(x^3+4xy^2+5y^3)=\dfrac{d}{dx}(1)"

Use the Chain Rule


"3x^2+4y^2+8xy\dfrac{dy}{dx}+15y^2\dfrac{dy}{dx}=0"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=-\dfrac{3x^2+4y^2}{8xy+15y^2}"

At "(2,-1)"


"\dfrac{dy}{dx}|_{(2,-1)}=-\dfrac{3(2)^2+4(-1)^2}{8(2)(-1)+15(-1)^2}=16"

"\dfrac{dy}{dx}|_{(2,-1)}=16"